OpenStudy (anonymous):

I roll a die repeatedly. Find the chance that the first 4 rolls all show different faces, and the 5th roll shows a face that has appeared before. unanswered

OpenStudy (anonymous):

You have 5 rolls. You want 4 different faces showing and a roll that shows any of the 4 different faces: r1 = 6/6 r2 = 5/6 r3 = 4/6 r4 = 3/6 r5 = 4/6 You want the prb of r1 and r2 and r3 and so on